OpenStudy (anonymous):

(10^25)-7 is divisible by a) 2 b)3 c)9 d) both 2 & 3

OpenStudy (anonymous):

b)3

OpenStudy (anonymous):

10^25=10000000000000000000000000 the number is not so important u know that the last digit will be 0 and is u substract 7 u will have tha last digit 3 si it is divisible by 3

OpenStudy (anonymous):

10-7=3 therefore 10^25-7 is divisible by 3.

OpenStudy (anonymous):

its not last digit but all preceeding digits are 9. therefore sum of 23 9s+3 is a multiple of 3.

OpenStudy (anonymous):

if u have 10 or 10^2 meaning 100 or 10^x it doesn't matter cause by substracting 7 u will have the last digit 3 and 3 is divisible only with 3

OpenStudy (anonymous):

133 for instance is not divisible by 3.

OpenStudy (anonymous):

yup, u are right

OpenStudy (shubhamsrg):

anyways,alternately, we know, 10^n = 3q + 1 and 7 = 3p + 1 subtracting both, 10^n - 7 = 3(q-p) = 3r ==> divisible by 3